OpenStudy (anonymous):

Solve for x. Show ALL work: -2x + 5 = 4x - 7

OpenStudy (anonymous):

-2x +5 = 4x -7 I then switch over the terms so I can add like terms: -2x - 4x = -7 - 5 After that I add the like terms -6x = -12 Finally, I divide -6 by both sides to get x by itself -6/-6x = -12/-6 = x = 2 Therefore: x = 2 :)

OpenStudy (anonymous):

thank you